OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, give the limit. 60, -10, five divided by three, negative five divided by eighteen, ...

OpenStudy (agent0smith):

You need to find the common ratio first - divide any term by the term right before it (use the first two terms since it's easiest to)

OpenStudy (anonymous):

i got -1/6

OpenStudy (agent0smith):

so the common ratio, r = -1/6. If |r| < 1, the sequence converges. Is |-1/6| < 1?

OpenStudy (anonymous):

The fraction looks like it will eventually be a small over large, so 0. That implies it should converge. The rest is finding the limit. Have you tried this reference for the topic? http://tutorial.math.lamar.edu/Classes/CalcII/ConvergenceOfSeries.aspx

OpenStudy (anonymous):

wait what

OpenStudy (agent0smith):

so the common ratio, r = -1/6. If |r| < 1, the sequence converges. Is |-1/6| < 1?

OpenStudy (anonymous):

1/6

OpenStudy (agent0smith):

Yes, but is |-1/6| < 1 ?

OpenStudy (anonymous):

no

OpenStudy (agent0smith):

Why not? 1/6 = 0.16 recurring.

OpenStudy (anonymous):

oh pellet yes because its aqbsolute value

OpenStudy (agent0smith):

Not because it's absolute value, but because 1/6th is less than 1.

OpenStudy (agent0smith):

So it converges, since |r| < 1

OpenStudy (agent0smith):

"If it converges, give the limit. " I guess they just want what the terms approach... I don't know if they want the value of the sum or not.

OpenStudy (anonymous):

Diverges Converges; 11100 Converges; 72 Converges; 0

OpenStudy (agent0smith):

The value of the sum: \[\Large \frac{ a_1 }{ 1-r }\]first term a1 and r, plug them in

OpenStudy (anonymous):

how do i get r though

OpenStudy (agent0smith):

You just found r :P

OpenStudy (agent0smith):

In the second post you gave it

OpenStudy (anonymous):

i got all -10

OpenStudy (agent0smith):

First term was 60, r=-1/6\[\Large \frac{ 60 }{ 1-\left( - \frac{ 1 }{ 6 }\right)}\]

OpenStudy (anonymous):

60.1666

OpenStudy (agent0smith):

Use a calculator, and make sure you use brackets.

OpenStudy (agent0smith):

http://www.wolframalpha.com/input/?i=+%5Cfrac%7B+60+%7D%7B+1-%5Cleft%28+-++%5Cfrac%7B+1+%7D%7B+6+%7D%5Cright%29%7D Looks like none of the answer choices are correct for the sum, so maybe they just meant the limit of the terms The terms will approach zero, so Converges; 0

OpenStudy (anonymous):

Oh man this was confusing lolol